OpenStudy (anonymous):

Fill in the blank to make the equivalent rational expressions. (-4)/(v+9)=(?)/(-v-9)

OpenStudy (anonymous):

\[\frac{ -4 }{ v+9 }=\frac{ ? }{ -v-9 }\] rearrange for ? So \[\frac{ 4v+36 }{ v+9 }= \frac{ 4(v+9) }{ v+9 }\] Can you simplify that?

hero (hero):

@sarahusher Why can't we just do this to begin with? \[\frac{-4}{v+9} = -\frac{4}{v+9} = \frac{4}{-(v + 9)}\]

hero (hero):

We can simplify that to figure out what's supposed to be in the question mark.

OpenStudy (anonymous):

@Hero I suppose you can, there are different ways of approaching it :)

OpenStudy (anonymous):

thank you